The following information is from the manufacturing budget and the budgeted financial statements of Fabor Fabrication.

  

Direct materials inventory, January 1 $ 68,000
Direct materials inventory, December 31 37,000
Direct materials budgeted for use during the year 190,000
Accounts payable to suppliers of materials, January 1 50,000
Accounts payable to suppliers of materials, December 31 79,000

a. Compute the budgeted amount for purchases of direct materials during the year.

b. Compute the budgeted amount for cash payments during the year to suppliers of materials.
Answer #1

a) Direct material purchase during the year = 190000+37000-68000 = 159000

b) Cash payment = 50000+159000-79000 = 130000